# Electron-Tunnelling-Noise Programmable Random Variate Accelerator for Monte Carlo Sampling

### v2

@@ 2 Programmable Random Variate Acceleration
1 Generate uniform [0, 1] random variate $u$
2 RETURN $x\leftarrow F^{-1}(u)$
@@
2       Generate uniform [0, 1] random variate $u$
3       Generate uniform [0, 1] random variate $x$
4       Set $T\leftarrow c\frac{f(x)}{g(x)}$
5
6until $uT\leq 1$
7RETURN $x$
@@ 4.2 Kernel Density Programmable Random Variate Accelerator for Programmable Univariate Distributions
1 $a=\frac{\sigma^{\prime}}{\sigma}$
2 $b=\mu^{\prime}-\mu a$
3 for $i=0$ to $N$ do
4       Read in random integer $x$ from the ADC
5       $\mathrm{Sample}=\frac{x+u}{2^{64}}$
6       $\mathrm{Samples}[i]=a*\mathrm{Sample}+b$
7
@@ 5 Noise Source Temperature Dependence
We placed the field programmable gate array development board and the Gaussian noise source inside a Binder MK56 thermal chamber. We set the thermal chamber to 0 C and allowed the temperature to reach equilibrium for 30 minutes. While waiting for the temperature to reach equilibrium we had the soft processor on the field programmable gate array consistently running a program to print raw analog-to-digital converter values. When equilibrium was reached we ran a program to have the processor sample and print $10^{6}$ raw analog-to-digital converter values. We repeated this for ten different temperatures between 0 C and 45 C with a 5 C temperature step.
@@
Figure 6: Mean and standard deviation of ten batches $10^{6}$ raw analog-to-digital converter samples over a temperature range of 0 C to 45 C with a 5 C temperature step. The blue data points and curves show the mean and standard deviation of the raw ADC values. The orange curve shows the standard deviation of the raw ADC values after 50 % of them have been randomly subtracted from the maximum ADC output value.
@@
Figure 7: Violin plot of kernel density estimates of $10^{6}$ ADC samples over a temperature range of 0 C to 45 C with a 5 C temperature step.

### v3

@@ 2 Programmable Random Variate Acceleration
Generate uniform [0, 1] random variate $u$ RETURN $x\leftarrow F^{-1}(u)$
@@
2       Generate uniform [0, 1] random variate $u$ Generate uniform [0, 1] random variate $x$ Set $T\leftarrow c\frac{f(x)}{g(x)}$
until $uT\leq 1$RETURN $x$
@@ 4.2 Kernel Density Programmable Random Variate Accelerator for Programmable Univariate Distributions
1 $a=\frac{\sigma^{\prime}}{\sigma}$ $b=\mu^{\prime}-\mu a$ for $i=0$ to $N$ do
2       Read in random integer $x$ from the ADC $\mathrm{Sample}=\frac{x+u}{2^{64}}$ $\mathrm{Samples}[i]=a*\mathrm{Sample}+b$
@@ 5 Noise Source Temperature Dependence
We placed the field programmable gate array development board and the Gaussian noise source inside a Binder MK56 thermal chamber. We set the thermal chamber to 0 ${}^{\circ}$ and allowed the temperature to reach equilibrium for 30 minutes. While waiting for the temperature to reach equilibrium we had the soft processor on the field programmable gate array consistently running a program to print raw analog-to-digital converter values. When equilibrium was reached we ran a program to have the processor sample and print $10^{6}$ raw analog-to-digital converter values. We repeated this for ten different temperatures between 0 ${}^{\circ}$ and 45 ${}^{\circ}$ with a 5 ${}^{\circ}$ temperature step.
@@
Figure 6: Mean and standard deviation of ten batches $10^{6}$ raw analog-to-digital converter samples over a temperature range of 0 ${}^{\circ}$ to 45 ${}^{\circ}$ with a 5 ${}^{\circ}$ temperature step. The blue data points and curves show the mean and standard deviation of the raw ADC values. The orange curve shows the standard deviation of the raw ADC values after 50 % of them have been randomly subtracted from the maximum ADC output value.
@@
Figure 7: Violin plot of kernel density estimates of $10^{6}$ ADC samples over a temperature range of 0 ${}^{\circ}$ to 45 ${}^{\circ}$ with a 5 ${}^{\circ}$ temperature step.